# Autonomous Bootstrapping of Quantum Dot Devices

### v1

@@ Abstract
Semiconductor quantum dots (QD) are a promising platform for multiple different qubit implementations, all of which are voltage-controlled by programmable gate electrodes. However, as the QD arrays grow in size and complexity, tuning procedures that can fully autonomously handle the increasing number of control parameters are becoming essential for enabling scalability. We propose a bootstrapping algorithm for initializing a depletion mode QD device in preparation for subsequent phases of tuning. During bootstrapping, the QD device functionality is validated, all gates are characterized, and the QD charge sensor is made operational. We demonstrate the bootstrapping protocol in conjunction with a coarse tuning module, showing that the combined algorithm can efficiently and reliably take a cooled-down QD device to a desired global state configuration in under minutes with a success rate of . Importantly, by following heuristic approaches to QD device initialization and combining the efficient ray-based measurement with the rapid radio-frequency reflectometry measurements, the proposed algorithm establishes a reference in terms of performance, reliability, and efficiency against which alternative algorithms can be benchmarked.
@@ I Introduction
Gate-defined quantum dots (QDs) in semiconductors present a promising avenue toward building high-fidelity and scalable quantum processors using existing industrial infrastructure [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, due to the gate-defined and gate-controlled nature of spin qubits based on QDs, effective tuning within the large-dimensional gate-voltage landscape is becoming a key limitation to qubit array operation [16]. Previous work has made progress in areas such as topology and charge state classification [17, 18, 19], navigation in higher-dimensional gate-voltage spaces using ray-based tuning [20, 14], and optimization of readout and qubit fidelity [21]. However, most algorithms assume that previous tuning steps, such as the definition of the QD(s) channel(s), calibration of the charge sensor(s), or identifying regions of interest within the coarse tuning regime, have been completed [18, 20, 19].
@@
In this work, we focus on the task of fully autonomous bootstrapping of the QD device, where “bootstrapping” is defined as a pretuning process that brings the device to an operational regime where the coarse tuning stage can be initiated [18, 16]. The proposed bootstrapping algorithm is designed to be initiated on a pristine QD device post-cooldown, for instance in the two-dimensional electron gas (2DEG) limit. While previous work proposed a machine learning (ML)-based approach to QD device initialization [22], our algorithm instead takes advantage of physics-based heuristics whenever possible. The reason we choose to rely on domain knowledge rather than ML methods is that, at present, even the best-performing ML models are limited in their interpretability, making it challenging to determine how the QD device should be adjusted in the case of failure. Moreover, low signal-to-noise ratio and other data imperfections can significantly affect the validity of ML predictions, causing unexpected failures of the tuning protocol. Finally, automating bootstrapping following heuristic approaches allows us to establish a reference in terms of performance, reliability, and efficiency against which alternative, possibly more ML-heavy algorithms can be benchmarked. Thus, the procedural flow of the proposed bootstrapping algorithm closely follows the typical manual tuning steps of a QD device. Moreover, to improve the tuning efficiency and reliability, our algorithm utilizes a proximal charge sensor [23], measured in radio-frequency (rf)-reflectometry, instead of transport measurements used in previous work [24].
@@
The bootstrapping process starts after cooling the device down to a few and involves determining the operational ranges, determining charging energies for each quantum dot (QD), and making the local charge sensing system operational. Figure 1(b) depicts the flow of the tuning protocol reported in this work. The procedural flow of the bootstrapping algorithm is shown in Fig. 1(c).
- Preparation.
@@
The first, preparatory stage [teal box in Fig. 1(c)] consists of setting the backbone (BB) gate to localize the operational region, followed by testing and setting the ohmic to a voltage that corresponds to the target current.
- Diagnostics and characterization.
@@
With the bias on the ohmic set, the protocol proceeds with diagnostics and characterization of the device finger gates [light blue box in Fig. 1(c)]. At this stage, the algorithm tests whether all relevant gates are functional and collects pinch-off and saturation voltages where applicable. Should any of the finger gates be flagged as non-functional, the bootstrapping process is terminated.
- Defining QDs channels.
@@
Defining the sensor QD and qubit QD channel [dark blue box in Fig. 1(c)] begins from a pinched-off current using the BB and right outer barrier (OBR). Then, the OBR is fine-tuned to the saturation point of the measured current.
- Setting up charge sensor.
@@
The final stage of bootstrapping consists of two parts. With the BB and OBR set, the algorithm first sets the sensor barrier gates (SBT and SBB) based on Coulomb oscillation in ray-based (rb) measurement and then the left outer barrier (OBL) in the qubit QD channel to its pinch-off [purple box in Fig. 1(c)]. Finally, the measurement system switches to rf-reflectometry, and the sensor plunger (SP) is calibrated [pink box in Fig. 1(c)].
@@
Once the bootstrapping process is completed, the tuning algorithm initiates the coarse tuning [18, 16]. The goal of coarse tuning is finding a range of gate voltages where the device is in a particular global configuration, i.e., a single-dot (SD) or double-dot (DD) state.
@@
We show that the proposed bootstrapping approach can be highly successful, with of test runs successfully preparing the device for coarse tuning. Given the rapid acquisition of data and its efficient analysis, our algorithm is capable of tuning the double-QD device within about minutes. Moreover, using the parameters gathered during bootstrapping, we show a success rate for coarse tuning, preparing the QD device for further calibration toward qubit operation.
@@
Finally, we demonstrate the cross-platform applicability of the ML model used in the coarse tuning module, which, thus far, has only been deployed on Si/SixGe1-x QD devices with an overlapping gate architecture [18, 17, 19, 25]. This work is, therefore, an important demonstration that an ML algorithm trained exclusively using synthetic data [26] can be successfully used to assess the state across various device designs.
@@
The remainder of the paper is organized as follows: In Sec. II we give an overview of the design of the bootstrapping protocol. The specifics of the algorithm and its configuration for all tests in this work are described in Sec. II.1 through Sec. II.4. The performance at the level of individual stages in bootstrapping is presented in Sec. III.1. The overall performance of the tuning algorithm in reaching the target global configuration is discussed in Sec. III.2. We conclude with a discussion of the potential modifications to further improve the proposed bootstrapping technique in Sec. IV.
@@ II Methodology: Experimental setup
Bootstrapping is initiated with a configuration file that defines the device architecture, identifies the quadrant that will be used in the experiment, defines the gates’ functionality (e.g., labeling gates as barriers or plungers), and specifies the device’s safety ranges. A scanning electron micrograph (SEM) of a gate-defined, depletion-mode GaAs/Al0.36Ga0.64As device used to demonstrate the bootstrapping protocol is shown in Fig. 1(a) [27]. The two-dimensional electron gas (2DEG) is formed 57 below the surface. The device, shown in Fig. 1(a), consists of four identical quadrants, each with tunable metallic electrodes for forming two QDs for hosting qubits and a larger sensor QD. Barrier gates [OBL, B, OBR, SBT, and SBB in Fig. 1(a) for outer left, middle, outer right, sensor top, and sensor bottom barrier, respectively] confine each of the QDs, while plunger gates [PL, PR, and SP in Fig. 1(a) for left, right, and sensor plunger, respectively] control their chemical potentials. The backbone gate and an appropriate separator gate [BB and OBR in Fig. 1(a)] are used to isolate the quadrant of interest. The separator gate is also used as a barrier for one of the qubit QDs as the designated gate was dysfunctional. Each quadrant has two ohmic contacts to the 2DEG which are also connected to an off-chip resonator. The ohmics are used to measure the current through the 2DEG. The resonator and sensor QD form an RLC circuit for rf-reflectometry measurements.
@@ II.1 Bootstrapping: Preparation
Once the device is cooled down, the preparatory stage begins [teal box in Fig. 1(c)]. The localization of the operational quadrant is achieved by setting the BB gate to the most negative safe voltage of the gate. Then, with all gates set to zero, the bias voltage is swept to test for ohmic behavior. If the expected linear relationship between the voltage bias and the measured current is observed, the ohmic is biased to the voltage where a small positive current of about can be observed. This amperage was determined during the manual prescreening to be both safe for the device, and sufficiently high for measurements using the experimental set-up described in Sec. II.
@@ II.2 Bootstrapping: Diagnostics and characterization
With the bias on the source ohmic set, the algorithm proceeds to assess the functionality of the quadrant’s finger gates, as shown in the light blue box in Fig. 1(c). This is achieved by applying a voltage to each gate, one at a time, and measuring the current through the device. Ideally, the voltage vs. current signal should have a sigmoidal shape. However, the device imperfections may lead to curves that more or less significantly violate this assumption, as can be seen in Fig. 2(b).
The analysis of the current vs. voltage signal serves several purposes. First, it is a diagnostic tool that verifies whether gates are functional and the device is suitable for tuning. Finger gates are classified as working if they produce a current vs. voltage signal with a sufficiently wide distribution, as quantified by the median absolute deviation (MAD). The MAD is compared to the threshold value that is expected to be higher than the experimental set-up noise levels determined from the device prescreening to be less than of the target current.
Secondly, all the relevant gate characteristics, such as pinch-off, saturation, and parking values, are determined as part of the bootstrapping process. To determine the pinch-off and saturation values, we consider the first () and the second () derivative of the current vs. voltage signal, respectively. Prior to analysis, the gradient of the signal is first smoothed with a Gaussian filter and then fit with an inverse cosine hyperbolic function of the form
where , , , and are fit parameters used for a post-hoc analysis. The location of the minimum-voltage prominent peak in the curve defines the pinch-off value. The saturation point value is defined at the location of the maximum-voltage negative prominent peak of the . The prominence assessment is used to identify stand-out features. For finding the pinch-off and saturation values, calculated prominence values are used to filter the peaks arising from the experimental set-up noise. Additionally, prominence analysis is used for finding and narrowing down the candidate list of optimal Coulomb oscillation peaks during the sensor QD tuning that uses a threshold value corresponding to of the target current, see Sec. II.4.
@@ II.3 Bootstrapping: Defining QDs channels
With all gates characterized, the bootstrapping algorithm proceeds to define the QDs channels [dark blue box in Fig. 1(c)]. The goal of this step is to adjust the BB and OBR gates to split the QD device quadrant into the sensor QD and qubit QD channels.
To find the combined voltages for BB and OBR that pinch off current, the algorithm utilizes the rb-measurement instead of the conventional 2D scans. The rb-approach involves measuring a collection of one-dimensional (1D) voltage sweeps initiated from a common starting point at a tilt from the central measurement axis defined as the straight line to . The tilt angles, set at six evenly spaced orientations in a range , are chosen to increase the likelihood of observing the current pinch-off. The algorithm picks the angle for which the pinch-off voltage is the lowest, decomposes the optimal pinch-off voltage into the BB and OBR voltage components, and then sets the BB gate based on this decomposition.
With BB calibrated, a full trace of OBR gate is measured and OBR is parked at the position of the peak in the with the highest current for sustaining transport through the QDs.
@@ II.4 Bootstrapping: Setting up charge sensor
Once the current channels are established, the protocol proceeds with tuning the outer barrier gates for both the sensor and qubit QDs [purple blue box in Fig. 1(c)]. With BB and OBR gates calibrated and the SP gate parked at 0 V, rb-measurement is conducted for SBT vs. SBB in a similar fashion as for BB vs. OBR.
@@
where is the peak height, is the peak’s full width at half maximum, and is an expected full width at half maximum for the device type used to test the bootstrapping algorithm [28]. The and are determined based on the prominence analysis described in Sec. II.2.
Once the most promising peak is determined, SP is put back into the open position, and the OBL is swept to localize qubit QDs. Again, if oscillations are observed, the qubit QD has been successfully confined, and thus, it is possible to split it with the B gate to form DD. We then proceed to turn on the rf-reflectometry setup for charge sensing. To achieve the best sensitivity, the plunger needs to be parked at the largest slope of the signal reflected off the sensor QD. This way small shifts in the DD capacitances due to electron transitions can be observed.
@@ II.5 Coarse tuning
Once the bootstrapping module is completed, the coarse tuning stage begins [18, 16]. The finger gate characteristics from bootstrapping are used to define the operational regime by limiting the range over which gates PL, B, and PR can be adjusted. The DSE-based tuning module is then initiated at voltages of the PL, B, and PR gates randomly selected from a uniform distribution between and for each respective gate, where . Proper scaling of the measurement scans is crucial for meaningful network analysis: scans that are too small may not contain enough features necessary for state classification, while scans that are too large may result in probability vectors that are not useful in the navigation phase. The size of the scans is determined based on the charging energies extracted during the diagnostics and characterization phase. The scan size is set to match roughly times the initial charging energies of individual QDs in the qubit QDs region of the device, as suggested in Ref. [25].
@@
This process is repeated until either the DD state is declared or the DSE module iteration limit is reached. Reaching the DSE module iteration limit triggers recalibration of the sensor QD dot, which is followed by reinitialization of the DSE module. By default, the iteration limit is set to . The reinitialization involves finding new starting values for PL, B, and PR. This is done by resampling the uniform distributions around saturation points for all three gates.
@@
Gate | Pinch-off [] | Saturation [] | Parking [] |

### v2

@@ Abstract
Semiconductor quantum dots (QDs) are a promising platform for multiple different qubit implementations, all of which are voltage controlled by programmable gate electrodes. However, as the QD arrays grow in size and complexity, tuning procedures that can fully autonomously handle the increasing number of control parameters are becoming essential for enabling scalability. We propose a bootstrapping algorithm for initializing a depletion-mode QD device in preparation for subsequent phases of tuning. During bootstrapping, the QD device functionality is validated, all gates are characterized, and the QD charge sensor is made operational. We demonstrate the bootstrapping protocol in conjunction with a coarse-tuning module, showing that the combined algorithm can efficiently and reliably take a cooled-down QD device to a desired global-state configuration in under min with a success rate of . Finally, by following heuristic approaches to QD device initialization and combining the efficient ray-based measurement with the rapid radio-frequency reflectometry measurements, the proposed algorithm establishes a reference in terms of performance, reliability, and efficiency against which alternative algorithms can be benchmarked.
@@ I Introduction
Gate-defined quantum dots (QDs) in semiconductors present a promising avenue toward building high-fidelity and scalable quantum processors using existing industrial infrastructure [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. However, due to the gate-defined and gate-controlled nature of spin qubits based on QDs, effective tuning within the large-dimensional gate-voltage landscape is becoming a key limitation to qubit array operation [16]. Previous work has made progress in areas such as topology and charge-state classification [17, 18, 19], navigation in higher-dimensional gate-voltage spaces using ray-based tuning [20, 14], and optimization of readout and qubit fidelity [21, 22]. However, most algorithms assume that previous tuning steps, such as the definition of the QD(s) channel(s), calibration of the charge sensor(s), or identifying regions of interest within the coarse-tuning regime, have been completed [18, 20, 19].
@@
In this work, we focus on the task of fully autonomous bootstrapping of the QD device, where “bootstrapping” is defined as a pretuning process that brings the device to an operational regime in which the coarse-tuning stage can be initiated [18, 16]. The proposed bootstrapping algorithm is designed to be initiated on a pristine QD device post-cooldown in the two-dimensional electron gas (2DEG) limit. While previous work has proposed a machine-learning- (ML) based approach to QD device initialization [23], our algorithm instead takes advantage of physics-based heuristics whenever possible. The reason we choose to rely on domain knowledge rather than ML methods is that, at present, even the best-performing ML models are limited in their interpretability, making it challenging to determine how the QD device should be adjusted in the case of failure. Moreover, low signal-to-noise ratio and other data imperfections can significantly affect the validity of ML predictions, causing unexpected failures of the tuning protocol. Finally, automating bootstrapping following heuristic approaches allows us to establish a reference in terms of performance, reliability, and efficiency against which alternative, possibly more ML-heavy, algorithms can be benchmarked. Thus, the procedural flow of the proposed bootstrapping algorithm closely follows the typical manual-tuning steps of a QD device. Moreover, to improve the tuning efficiency and reliability, our algorithm utilizes a proximal charge sensor [24], measured in radio-frequency (rf) reflectometry, instead of transport measurements used in previous work [25].
@@
The bootstrapping process starts after cooling the device down to a few millikelvins and involves determining the operational ranges, determining charging energies for each quantum dot (QD), and making the local charge-sensing system operational. In Fig. 1(b), we depict the flow of the tuning protocol reported in this work. The procedural flow of the bootstrapping algorithm is shown in Fig. 1(c).
@@
Preparation. The first, preparatory stage [teal box in Fig. 1(c)] consists of setting the backbone (BB) gate to localize the operational region, followed by testing and setting the Ohmic to a voltage that corresponds to the target current.
@@
Diagnostics and characterization. With the bias on the Ohmic set, the protocol proceeds with diagnostics and characterization of the device finger gates [light blue box in Fig. 1(c)]. At this stage, the algorithm tests whether all relevant gates are functional and collects pinch-off and saturation voltages where applicable. Should any of the finger gates be flagged as non-functional, the bootstrapping process is terminated.
@@
Defining QDs channels. Defining the sensor-QD and qubit-QD channel [dark blue box in Fig. 1(c)] begins from a pinched-off current using the BB and right outer barrier (OBR). Then, the OBR is fine tuned to the saturation point of the measured current.
@@
Setting up charge sensor. The final stage of bootstrapping consists of two parts. With the BB and OBR set, the algorithm first sets the sensor barrier gates (SBT and SBB) based on Coulomb oscillation in ray-based (rb) measurement and then the left outer barrier (OBL) in the qubit-QD channel to its pinch-off [purple box in Fig. 1(c)]. Finally, the measurement system switches to rf reflectometry, and the sensor plunger (SP) is calibrated [pink box in Fig. 1(c)].
@@
Once the bootstrapping process is completed, the tuning algorithm initiates the coarse tuning [18, 16]. The goal of coarse tuning is to find a range of gate voltages where the device is in a particular global configuration, i.e., a single-dot (SD) or double-dot (DD) state.
@@
We show that the proposed bootstrapping approach can be highly successful, with of test runs successfully preparing the device for coarse tuning. Given the rapid acquisition of data and its efficient analysis, our algorithm is capable of tuning the double-QD device within about min. Moreover, using the parameters gathered during bootstrapping, we show a success rate for coarse tuning, preparing the QD device for further calibration toward qubit operation.
@@
Finally, we demonstrate the cross-platform applicability of the ML model used in the coarse-tuning module, which, thus far, has only been deployed on Si/SixGe1-x QD devices with an overlapping gate architecture [18, 17, 19, 26]. This work is, therefore, also a demonstration that an ML algorithm trained exclusively using synthetic data [27] can be successfully used to assess the state across various device designs.
@@
The remainder of the paper is organized as follows: In Sec. II, we give an overview of the design of the bootstrapping protocol. The specifics of the algorithm and its configuration for all tests in this work are described in Secs. II.1– II.4. The performance at the level of individual stages in bootstrapping is presented in Sec. III.1. The overall performance of the tuning algorithm in reaching the target global configuration is discussed in Sec. III.2. We conclude with a discussion of the potential modifications to further improve the proposed bootstrapping technique in Sec. IV.
@@ II Methodology: Experimental setup
Bootstrapping is initiated with a configuration file that defines the device architecture, identifies the quadrant that will be used in the experiment, defines the functionality of the gates (e.g., labeling gates as barriers or plungers), and specifies the safety ranges of the device. A scanning electron micrograph (SEM) of a gate-defined, depletion-mode GaAs/Al0.36Ga0.64As device used to demonstrate the bootstrapping protocol is shown in Fig. 1(a) [28]. The two-dimensional electron gas (2DEG) is formed 57 below the surface. The device, shown in Fig. 1(a), consists of four identical quadrants, each with tunable metallic electrodes for forming two QDs for hosting qubits and a larger sensor QD. Barrier gates [OBL, B, OBR, SBT, and SBB in Fig. 1(a) for the outer left, middle, outer right, sensor top, and sensor bottom barrier, respectively] confine each of the QDs, while plunger gates [PL, PR, and SP in Fig. 1(a) for the left, right, and sensor plunger, respectively] control their chemical potentials. The backbone gate and an appropriate separator gate [BB and OBR in Fig. 1(a)] are used to isolate the quadrant of interest. The separator gate is also used as a barrier for one of the qubit QDs as the designated gate was dysfunctional. Each quadrant has two Ohmic contacts to the 2DEG which are also connected to an off-chip resonator. The Ohmics are used to measure the current through the 2DEG. The resonator and sensor QD form an RLC circuit for rf-reflectometry measurements.
@@ II.1 Bootstrapping: Preparation
Once the device is cooled down, the preparatory stage begins [teal box in Fig. 1(c)]. The localization of the operational quadrant is achieved by setting the BB gate to the most negative safe voltage of the gate. Then, with all gates set to zero, the bias voltage is swept to test for Ohmic behavior. If the expected linear relationship between the voltage bias and the measured current is observed, the Ohmic is biased to the voltage where a small positive current of about can be observed. This amperage was determined during the manual prescreening to be both safe for the device, and sufficiently high for measurements using the experimental setup described in Sec. II.
@@ II.2 Bootstrapping: Diagnostics and characterization
With the bias on the source Ohmic set, the algorithm proceeds to assess the functionality of the finger gates in the quadrant, as shown in the light blue box in Fig. 1(c). This is achieved by applying a voltage to each gate, one at a time, and measuring the current through the device. Ideally, the voltage versus current signal should have a sigmoidal shape. However, the device imperfections may lead to curves that more or less significantly violate this assumption, as can be seen in Fig. 2(b).
The analysis of the current versus voltage signal serves several purposes. First, it is a diagnostic tool that verifies whether gates are functional and the device is suitable for tuning. Finger gates are classified as working if they produce a current versus voltage signal with a sufficiently wide distribution, as quantified by the median absolute deviation (MAD). The MAD is compared to the threshold value that is expected to be higher than the experimental setup noise levels determined from the device prescreening to be less than of the target current.
Second, all the relevant gate characteristics, such as pinch-off, saturation, and parking values, are determined as part of the bootstrapping process. To determine the pinch-off and saturation values, we consider the first () and the second () derivative of the current versus voltage signal, respectively. Prior to analysis, the gradient of the signal is first smoothed with a Gaussian filter and then fitted with an inverse cosine hyperbolic function of the form
where , , , and are fit parameters used for a post-hoc analysis [29]. The location of the minimum-voltage prominent peak in the curve defines the pinch-off value. The saturation point value is defined at the location of the maximum-voltage negative prominent peak of the . The prominence assessment is used to identify stand-out features. For finding the pinch-off and saturation values, calculated prominence values are used to filter the peaks arising from the experimental setup noise. Additionally, prominence analysis is used for finding and narrowing down the candidate list of optimal Coulomb oscillation peaks during the sensor-QD tuning that uses a threshold value corresponding to of the target current (see Sec. II.4).
@@ II.3 Bootstrapping: Defining QDs channels
With all gates characterized, the bootstrapping algorithm proceeds to define the QDs channels [dark blue box in Fig. 1(c)]. The goal of this step is to adjust the BB and OBR gates to split the QD device quadrant into the sensor-QD and qubit-QD channels.
To find the combined voltages for BB and OBR that pinch off current, the algorithm utilizes the rb measurement instead of the conventional 2D scans. The rb approach involves measuring a collection of one-dimensional (1D) voltage sweeps initiated from a common starting point at a tilt from the central measurement axis, defined as the straight line to . The tilt angles, set at seven evenly spaced orientations in a range between and , are chosen to increase the likelihood of observing the current pinch-off. The algorithm picks the angle for which the pinch-off voltage is the lowest, decomposes the optimal pinch-off voltage into the BB and OBR voltage components, and then sets the BB gate based on this decomposition.
With BB calibrated, a full trace of the OBR gate is measured and OBR is parked at the position of the peak in the with the highest current for sustaining transport through the QDs.
@@ II.4 Bootstrapping: Setting up charge sensor
Once the current channels are established, the protocol proceeds with tuning the outer barrier gates for both the sensor and qubit QDs [purple blue box in Fig. 1(c)]. With the BB and OBR gates calibrated and the SP gate parked at 0 V, rb measurement is conducted for SBT versus SBB in a similar fashion as for BB versus OBR.
@@
where is the peak height, is the full width at half maximum of the peak, and is an expected full width at half maximum for the device type used to test the bootstrapping algorithm [30]. The and are determined based on the prominence analysis described in Sec. II.2.
Once the most promising peak is determined, SP is put back into the open position, and the OBL is swept to localize qubit QDs. Again, if oscillations are observed, the qubit QD has been successfully confined and, thus, it is possible to split it with the B gate to form DD. We then proceed to turn on the rf-reflectometry setup for charge sensing. To achieve the best sensitivity, the plunger needs to be parked at the largest slope of the signal reflected off the sensor QD. In this way, small shifts in the DD capacitances due to electron transitions can be observed.
@@ II.5 Coarse tuning
Once the bootstrapping module is completed, the coarse-tuning stage begins [18, 16]. The finger-gate characteristics from bootstrapping are used to define the operational regime by limiting the range over which gates PL, B, and PR can be adjusted. The DSE-based tuning module is then initiated at voltages of the PL, B, and PR gates randomly selected from a uniform distribution between and for each respective gate, where . Proper scaling of the measurement scans is crucial for meaningful network analysis: scans that are too small may not contain enough features necessary for state classification, while scans that are too large may result in probability vectors that are not useful in the navigation phase. The size of the scans is determined based on the charging energies extracted during the diagnostics and characterization phase. The scan size is set to match roughly times the initial charging energies of individual QDs in the qubit-QDs region of the device, as suggested in Ref. [26].
@@
This process is repeated until either the DD state is declared or the DSE module iteration limit is reached. Reaching the DSE module iteration limit triggers recalibration of the sensor QD, which is followed by reinitialization of the DSE module. By default, the iteration limit is set to . The reinitialization involves finding new starting values for PL, B, and PR. This is done by resampling the uniform distributions around saturation points for all three gates.
@@
Gate | Pinch-off () | Saturation () | Parking () |